# Quantum querying based on multicontrolled Toffoli gates for causal Feynman loop configurations and directed acyclic graphs

### v1

@@ Abstract
Quantum algorithms are a promising framework for a proper treatment of Feynman loop integrals due to the existence of a manifestly causal representation scenario. Particularly, unfolding causal configurations of multiloop Feynman diagrams is understood as querying directed acyclic graph (DAG) configurations of undirected graphs in graph theory. In this paper we present a quantum algorithm for querying causality of multiloop Feynman diagrams using an ingenious change in the logic of the design of the oracle operator. The construction of the quantum oracle is surprisingly based exclusively on multicontrolled Toffoli gates and XNOT gates. The efficiency of the algorithm is evaluated performing a comparison with a quantum algorithm based on binary clauses. Additionally, we explicitly analise several three-, four- and five-eloop topologies, which have not been previously explored due to their higher complexity.
@@ 1 Introduction
Quantum computing Feynman:1981tf is currently an appealing approach with great potential to tackle complex problems in many fields where the quantum principles of superposition and entanglement suppose an advantage. In the context of high-energy physics Humble:2022klb ; Rodrigo:2024say recent applications of quantum algorithms include lattice gauge theories Jordan:2011ne ; Banuls:2019bmf ; Zohar:2015hwa ; Byrnes:2005qx , the speed up of track reconstruction Magano:2021jzd ; Duckett:2022ccc ; Schwagerl:2023elf and jet clustering algorithms Wei:2019rqy ; Pires:2020urc ; deLejarza:2022bwc ; deLejarza:2022vhe , analysing the formation of jets in a medium Barata:2021yri ; Barata:2022wim ; Barata:2023clv , simulation of parton showers Bauer:2019qxa ; Bauer:2021gup ; Bepari:2020xqi ; Williams:2021lvr , determination of parton densities Perez-Salinas:2020nem ; Cruz-Martinez:2023vgs , helicity amplitudes Bepari:2020xqi and the colour algebra Chawdhry:2023jks of elementary processes, heavy-ion collisions deJong:2020tvx , quantum machine learning Guan:2020bdl ; Wu:2020cye ; Trenti:2020ceh , and quantum integrators Herbert:2021xgs ; Agliardi:2022ghn ; deLejarza:2023qxk , including their application to Feynman loop integrals deLejarza:2024pgk . In this work, we delve into the quantum determination of the causal structure of multiloop Feynman diagrams previously analysed in  Ramirez-Uribe:2021ubp ; Clemente:2022nll .
@@
Improving the accuracy of theoretical predictions for high-energy colliders is crucial because of the high demands this field will face Strategy:2019vxc , in particular during the current Run 3 of the CERN’s Large Hadron Collider (LHC), the planned high-luminosity LHC stage Gianotti:2002xx , and future collider projects Abada:2019lih ; Djouadi:2007ik ; Roloff:2018dqu ; CEPCStudyGroup:2018ghi . As far as perturbative Quantum Field Theory is concerned, one of the main needs is the study and appropriate treatment of vacuum quantum fluctuations associated to multiloop Feynman diagrams. In accordance with this need, there has been a remarkable effort in the field to provide powerful frameworks to meet this challenge Heinrich:2020ybq .
Aiming for improving the efficiency of multiloop-level computations with a proper treatment of singularities, we work in the novel framework of the Loop-Tree Duality (LTD) Catani:2008xa ; Bierenbaum:2010cy ; Bierenbaum:2012th ; Tomboulis:2017rvd ; Runkel:2019yrs ; Capatti:2019ypt , which opens any loop diagram into a sum of connected trees. The LTD framework has been extensively studied Buchta:2014dfa ; Hernandez-Pinto:2015ysa ; Jurado:2017xut ; Driencourt-Mangin:2019yhu ; Aguilera-Verdugo:2019kbz and several applications have been developed Buchta:2015wna ; Sborlini:2016gbr ; Sborlini:2016hat ; Driencourt-Mangin:2017gop ; Driencourt-Mangin:2019aix ; Capatti:2019edf ; Plenter:2019jyj ; Prisco:2020kyb ; Plenter:2020lop ; Runkel:2019zbm . The latest achievements through the LTD Verdugo:2020kzh ; snowmass2020 ; Aguilera-Verdugo:2020kzc ; Aguilera-Verdugo:2020nrp ; Ramirez-Uribe:2020hes ; Sborlini:2021owe ; TorresBobadilla:2021ivx ; TorresBobadilla:2021dkq ; Aguilera-Verdugo:2021nrn ; Ramirez-Uribe:2021ubp ; Rios-Sanchez:2024xtv ; Clemente:2022nll are based on its most distinctive and exceptional feature: the existence of a manifestly causal representation and its connection with directed acyclic graph (DAG) configurations in graph theory. The emergence of noncausal singularities in the customary Feynman representation imply considerable numerical instabilities whereas with the causal LTD representation scenario, noncausal singularities are explicitly absent and provide the advantage of significantly more stable integrands, among other appealing properties Aguilera-Verdugo:2020kzc ; Ramirez-Uribe:2020hes .
The fact that the two on-shell states of a Feynman propagator can be naturally encoded in a qubit makes quantum computing a potential efficient methodology to treat Feynman loop integrals. In Ref. Ramirez-Uribe:2021ubp , a quantum algorithm which succeeded to unfold the causal configurations of multiloop Feynman diagrams was proposed. A similar study based on a variational quantum eigensolver approach has been presented in Ref. Clemente:2022nll . In this paper we present an enhanced quantum algorithm for querying causality of multiloop Feynman diagrams with a significant reduction of qubit requirement and quantum circuit depth reduction EPpatent . The improvement achieved allows us to explore multiloop topologies of higher complexity that were previously unreachable.
@@ 2 Causality from the loop-tree duality
Generic loop integrals and scattering amplitudes, with $n$ propagators and $P$ external legs are defined, in the Feynman representation, as integrals in the Minkowski space of the $L$ loop momenta
${\cal A}_{F}^{(L)}=\int_{\ell_{1}\ldots\ell_{L}}{\cal N}\big{(}\{\ell_{s}\}_{L},\{p_{j}\}_{P}\big{)}\prod_{i=1}^{n}G_{F}(q_{i})~{},$ (1)
where the momenta flowing through each Feynman propagator, $q_{i}$, is a linear combination of the loop and external momenta, and the numerator ${\cal N}$ is given by the specific theory. The integration measure in dimensional regularisation Bollini:1972ui ; tHooft:1972tcz reads $\int_{\ell_{s}}=-\imath\mu^{4-d}\int{\rm d}^{d}\ell_{s}/(2\pi)^{d}$, with $d$ the number of space-time dimensions and $\mu$ an arbitrary energy scale. Feynman propagators are written conveniently as
$G_{F}(q_{i})=\frac{1}{\big{(}q_{i,0}-q_{i,0}^{(+)}\big{)}\big{(}q_{i,0}+q_{i,0}^{(+)}\big{)}}~{},$ (2)
where $q_{i,0}^{(+)}=\sqrt{\mathbf{q}_{i}^{2}+m_{i}^{2}-\imath 0}$, $q_{i,0}$ and $\mathbf{q}_{i}$ are respectively the temporal and spacial components of $q_{i}$, and $m_{i}$ is the mass of the propagating particle. The infinitesimal factor $\imath 0$ is the usual complex prescription of a Feynman propagator.
Singularities in the integrand of Eq. (1) arise when Feynman propagators are set on shell, explicitly when the energy component of $q_{i,0}$ takes either value of ±$\pm q_{i,0}^{(+)}$ in Eq. (2). An important fact to take into account, is that not all potential singular configurations of the integrand in Eq. (1) lead to physical singularities of the integral. In the case of the Feynman representation of loop integrals, noncausal singularities are unavoidable. Regarding the LTD framework, its more remarkable feature is the existence of a manifestly causal representation. The direct LTD representation of Eq. (1) is computed by applying the Cauchy’s residue theorem through the evaluation of nested residues Verdugo:2020kzh . To obtain the causal dual representation we sum over all the nested residues, explicitly cancelling all the noncausal contributions. Furthermore, a more suitable dual causal representation is found by cleverly reinterpreting it in terms of entangled thresholds Aguilera-Verdugo:2020kzc ; Ramirez-Uribe:2020hes . The analytical reconstruction is achieved by matching all combinations of $n-L$ thresholds that are causally compatible to each other, leading to the LTD causal representation,
${\cal A}_{D}^{(L)}=\int_{\vec{\ell}_{1}\ldots\vec{\ell}_{L}}\frac{1}{x_{n}}\sum_{\sigma\in\Sigma}\frac{{\cal N}_{\sigma(i_{1},\ldots,i_{n-L})}}{\lambda_{\sigma(i_{1})}^{h_{\sigma(i_{1})}}\cdots\lambda_{\sigma(i_{n-L})}^{h_{\sigma(i_{n-L})}}}+(\lambda_{p}^{+}\leftrightarrow\lambda_{p}^{-})~{},$ (3)
with $x_{n}=\prod_{i=1}^{n}2q_{i,0}^{(+)}$ and $h_{\sigma(i)}=\pm$. The Feynman propagators from Eq. (1) are substituted in Eq. (3) by causal propagators $1/\lambda_{p}^{\pm}$, with
@@
where $p$ is a partition of the on-shell energies, and $k_{p,0}$ is a linear combination of the external momenta energy components. Given the sign of $k_{p,0}$, either $\lambda_{p}^{-}$ or $\lambda_{p}^{+}$ becomes singular after all the propagators in $p$ are set on shell. The combinations of entangled causal propagators represent causal thresholds that can occur simultaneously which are collected in the set $\Sigma$.
A crucial step in the procedure for obtaining the causal representation shown in Eq. (3), is the identification of diagram configurations admitting compatibility among causal thresholds (causal singular configurations), which can be seen as querying multiple solutions over an unstructured database Boyer:1996zf . This exercise is equivalent to select all the DAG configurations of an undirected graph in graph theory. In the following section we present an efficient quantum algorithm to address the identification of causal singular configurations of multiloop Feynman diagrams.
@@ 3 Quantum causal querying
The proposed algorithm recovers the central ideas from Ref. Ramirez-Uribe:2021ubp for the applicability of amplitude amplification or Grover’s quantum algorithm Grover:1997fa ; Boyer:1996zf and considers a generalised description of multiloop topologies by treating them as directed graphs made of edges and eloops TorresBobadilla:2021ivx ; Sborlini:2021owe . An edge is defined as the union of multiple propagators connecting two interaction vertices. Once propagators are substituted by edges, the remaining loops of the graph define the number of eloops, which is smaller than the number of loops of the original Feynman diagram. The convenience of working with eloops in a causality scenario is given by the fact that the only feasible causal configurations are those for which the momentum flows of all propagators belonging to each edge are aligned in the same direction.
The distinguishing idea introduced in this paper is a clever change of logic in the construction of the oracle operator. The tagging of the unidirectional cycles representing eloops is done through the application of only multicontrolled Toffoli gates and XNOT gates, having as a core difference with Ref. Ramirez-Uribe:2021ubp the absence of binary clauses comparing adjacent edges. To evaluate the efficiency of the proposal, we will focus on the number of qubits required for the algorithm implementation and on the quantum circuit depth. Before describing the algorithm, we review some relevant concepts for a better understanding of the work developed.
Regarding the applicability of Grover’s quantum algorithm, the mixing angle that measures the ratio of states to be queried is defined as $\theta=\arcsin\sqrt{r/N}$, with $r$ the number of winning states and $N=2^{n}$ the total number of possible states. This value is crucial to determine the feasibility of the methodology, in fact, the optimal number of iterations, $t$, to obtain a feasible solution from Grover’s algorithm is achieved when $\sin^{2}\theta_{t}\sim 1$ with $\theta_{t}=(2t+1)\,\theta$. It is worth mentioning that the most effective implementation for the majority of the multiloop topologies analysed is achieved with a single iteration, $t=1$, when the initial mixing angle is $\theta\sim\pi/6$, or correspondingly, when $r/N\sim 1/4$.
@@
It has been shown from classical Aguilera-Verdugo:2020kzc ; Ramirez-Uribe:2020hes and quantum Ramirez-Uribe:2021ubp computations that the number of causal configurations of the multiloop topologies depicted in Fig. 1 is about one half of the total number of possible states. In order to achieve a succesful applicability of the quantum algorithm, there are two actions that can be undertaken. As a first option, we can halve the number of causal solutions to be queried by considering configurations with an edge fixed in one state. This can be done by either fixing a qubit in the oracle’s marker Ramirez-Uribe:2021ubp or excluding it from the circuit assuming that the qubit state is given. Halving the solutions to be queried is viable due to the fact that the objects of study are Feynman loop diagrams, which have the feature that given a causal solution the mirror state with all the momentum flows reversed is also a causal solution. A second adjustment, just in case the scenario requires it, is to introduce an extra qubit Nielsen:2012yss to increase the number of total states without increasing the number of solutions.
Moving on to the implementation of the quantum algorithm, three registers are needed. The first register, $|{e}\rangle$, encodes the states of all edges. We define the Boolean functions
@@
where $s_{j}$ labels the set of edges $e_{i}$ depending on the same linear combination of loop momenta. The bar in $\bar{s}_{j}$ denotes that the qubit states in this set are inverted with XNOT gates, $|{\bar{e}_{i}}\rangle=X|{e_{i}}\rangle$. The number of edges in the set $s_{j}$ is denoted by $n_{s_{j}}$ and the total number of edges in the corresponding multiloop topology is $n=\sum_{j}n_{s_{j}}$. The second register required, $|{a}\rangle$, collects the causal conditions in eloop clauses. The last register, $|{out}\rangle$, stores the validation of the causal solutions in what is known as the oracle’s marker.
The algorithm takes as a first step the initialisation of all the quantum registers. The register encoding the edges’ states, $|{e}\rangle$, is initialised in a uniform superposition by applying Hadamard gates to each of its $n$ qubits, $|{e}\rangle=H^{\otimes n}|{0}\rangle$. For the eloop clauses, the $m$ qubits in the register $|{a}\rangle$ are initialised to $|{1}\rangle$ by the application of Pauli-X gates, $|{a}\rangle=X^{\otimes m}|{0}\rangle$. The oracle’s marker qubit, $|{out}\rangle$, is initialised to one of the Bell states by applying $|{out}\rangle=H(X|{0}\rangle)=H|{1}\rangle\equiv|{-}\rangle$.
Following to the construction of the oracle, it is important to recall that causal configurations are identified as DAG configurations, therefore, the eloop clauses are established to suppress directed cyclic configurations. The eloop clauses are implemented through the application of multicontrolled Toffoli gates and XNOT gates that validate whether a configuration is cyclic or acyclic. To provide a convenient notation, we define the multicontrolled Toffoli operation based on Boolean clauses by
@@
where the capital $S_{j}$ contains the set of edges generating the cyclic configuration involving the subloop $j$. Both conditions in Eq. (6) select the same subloop but with opposite edge directions. The composite condition in Eq. (7) embraces both directions of the given subloop associated to $S_{j}$. Let us recall that the eloop clause register $|{a}\rangle$ is initialised to $|{1}\rangle$, therefore, applying one of the conditions in Eqs. (6) and (7) unmarks a noncausal configuration from $|{a}\rangle$.
The number of eloop clauses depends on the number of subloops needed to validate. An important aspect to consider relating eloop clauses is whether or not the set $S_{j}$ involves a fixed qubit, that is, a qubit in a definite state $|{0}\rangle$ or $|{1}\rangle$. If the eloop clause condition does not involve a fixed qubit, we apply Eq. (7), which considers two multicontrolled Toffoli gates to validate both directions. The validation of both directions is stored in a single qubit $|{a_{i}}\rangle$, due to the fact that both configurations are mutually exclusive, i.e., applying one multicontrolled Toffoli gate do not overlap over the second one. In the case that the eloop clause involves a fixed qubit, the condition to prove involves only one direction, explicitly, only one of the two Boolean conditions given in Eq. (6) is tested.
@@
$U_{w}|{e}\rangle|{a}\rangle|{out}\rangle=(-1)^{f(a,e)}|{e}\rangle|{a}\rangle|{out}\rangle~{},$ (8)
where $f(a,e)=\left(\bigwedge_{i}a_{i}\right)\wedge e_{0}$, and $e_{0}$ is the edge whose state is fixed. The procedure continues by performing the oracle operations in inverse order to restore every qubit but $|{out}\rangle$ to its initial state. At this point we are ready to apply the diffusion operator, $U_{s}$, to the register $|{e}\rangle$. The diffusion operator implemented is the one described in the IBM Qiskit website  [^1] http://qiskit.org/.
The quantum algorithm presented in this paper has the advantage that the quantum circuit depth depends just on the number of eloop clauses and not on the number of edges per set, whereas the quantum circuit depth of the quantum algorithm using binary clauses Ramirez-Uribe:2021ubp increases when additional edges per set are considered.
Let us recall that the quantum circuit depth Depth:5769 is defined as the integer number of gates that need to be executed along the longest path of the circuit from input to output, moving forward in time along qubit wires. The input is understood as the preparation of the qubits and the output as the measurement gate. The quantum circuit depth is computed through the Qiskit routine, QuantumCircuit.depth() [^2] https://qiskit-test.readthedocs.io. An important feature of the quantum circuit depth is the impact of quantum gates acting on no common qubits, allowing to perform them at the same time-step. On the other hand, if the gates act on at least one common qubit they have to be applied in different time-steps, increasing by one unit the quantum circuit depth. In the forthcoming sections we explore in a detailed way the performance of the algorithm by analysing the number of qubits needed for the implementation and the corresponding theoretical quantum circuit depth.
@@
Figure 2: Top left: Two eloop topology with two edges per set. Top right: Improved quantum circuit with an oracle based on multicontrolled Toffoli gates. Bottom: Implementation of the quantum algorithm presented in Ref Ramirez-Uribe:2021ubp using CNOT gates for the binary clauses and multicontrolled Toffoli gates for the registers $|{a}\rangle$ and $|{out}\rangle$.
@@
For the oracle’s marker, another multicontrolled Toffoli gate is used. For this case we achieve an initial mixing angle of $36.8^{\circ}$ (see Table. 1) by fixing the qubit $|{e_{0}}\rangle$. The multicontrolled Toffoli gate implementing the oracle’s marker takes as control the qubits in the register $|{a}\rangle$ and the qubit $|{e_{0}}\rangle$, and as target the qubit $|{out}\rangle$. After applying the oracle operations in reverse order, we apply the diffusion operator to the register $|{e}\rangle$. Table. 1 shows the results of the two quantum circuits in Fig. (2), exhibiting a clear improvement, going from $16$ to $10$ total number of qubits for implementation, and from $22$ to $14$ quantum depth.
Fig. eloops (edges) Qubits $\begin{subarray}{c}\text{Quantum}\\ \text{Depth}\end{subarray}$ $\theta$ $\sin^{2}(\theta_{t})$ $\begin{subarray}{c}\text{Causal}\\ \text{states}\end{subarray}$
2 two ($6$) $\phantom{0}10~{}|~{}16$ $14~{}|~{}22$ $36.8^{\circ}$ $0.87$ $23$
Table 1: Quantum circuit requirements of the quantum algorithms used to analyse the two-eloop topology in Fig. 2. The first number in the second and third columns is from the quantum algorithm proposed in this work, whereas the second number corresponds to Ref. Ramirez-Uribe:2021ubp . The columns with a single entry have the same value for both algorithms. The number in the last column is the number of causal configurations.

### v2

@@ Abstract
Quantum algorithms are a promising framework for unfolding the causal configurations of multiloop Feynman diagrams, which is equivalent to querying the directed acyclic graph (DAG) configurations of undirected graphs in graph theory. In this paper, we present a quantum algorithm for querying in both types of applications, using a systematic and sparing logic in the design of an oracle operator. The construction of the quantum oracle is based exclusively on multicontrolled Toffoli (MCX) gates and quantum NOT (Pauli-$X$) gates. The efficiency of the algorithm is evaluated by comparison with a quantum algorithm based on binary clauses. Furthermore, we analyse the impact of traspilation and introduce an appropriate metric to assess the complexity of the algorithm, the quantum circuit area. We explicitly analyse three-, four- and five-eloop topologies, which have not previously been explored due to their higher complexity and the current limitations of quantum simulators.
@@ 1 Introduction
Quantum computing [1] is currently an appealing approach with great potential to tackle complex problems in many fields where exploiting the quantum principles of superposition and entanglement could suppose an advantage. In the context of high-energy physics [2, 3] recent applications of quantum algorithms include lattice gauge theories [4, 5, 6, 7], track reconstruction [8, 9, 10], jet clustering algorithms [11, 12, 13, 14], analysing the formation of jets in a medium [15, 16, 17], simulation of parton showers [18, 19, 20, 21], determination of parton densities [22, 23], helicity amplitudes [20] and the colour algebra [24] of elementary processes, heavy-ion collisions [25], quantum machine learning [26, 27, 28], and quantum integrators [29, 30, 31], including their application to loop Feynman integrals and decay rates at next-to-leading order [32, 33]. In this work, we delve into the quantum determination of the causal structure of multiloop Feynman diagrams previously analysed in  [34, 35].
@@
Improving the accuracy of theoretical predictions for high-energy colliders is crucial because of the high demands this field will face [36], in particular during the current Run 3 of the CERN’s Large Hadron Collider (LHC), the planned high-luminosity LHC stage [37], and future collider projects [38, 39, 40, 41]. As far as perturbative Quantum Field Theory is concerned, one of the main needs is the study and appropriate treatment of the quantum fluctuations associated to multiloop Feynman diagrams. In accordance with this need, there has been a remarkable effort in the field to provide powerful frameworks to meet this challenge [42].
Aiming for improving the efficiency of multiloop-level computations with a proper treatment of singularities, we work in the framework of the Loop-Tree Duality (LTD) [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. In particular, the latest achievements in LTD are based on its most distinctive feature: the existence of a manifestly causal representation and its connection with directed acyclic graph (DAG) configurations in graph theory [64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]. A comprehensive review of classical algorithms in graph theory is presented in Ref. [76].
The emergence of noncausal singularities in the customary Feynman representation of loop integrals implies considerable numerical instabilities, whereas in the LTD representation, noncausal singularities are explicitly absent, providing the advantage of significantly more stable integrands [66, 68]. The fact that a Feynman propagator has two on-shell states, which are therefore naturally encoded in a qubit, makes quantum computing a potentially efficient way of dealing with some aspects of multiloop Feynman integrals. In Ref. [34], an amplitude amplification or Grover’s based [77, 78] quantum algorithm was proposed to unfold the causal configurations of multiloop Feynman diagrams. A similar study based on a variational quantum eigensolver (VQE) approach was presented in Ref. [35]. The quantum algorithm of Ref. [34] is quite expensive in terms of quantum resources, while the VQE approach [35] requires fewer quantum resources but longer runs.
In this paper, we present an enhanced quantum algorithm for querying causality of multiloop Feynman diagrams with a significant reduction in qubit requirements and quantum circuit area through a more systematic and efficient logic implementation for identifying directed acyclic graphs [79]. The improvement achieved allows us to explore multiloop topologies of higher complexity that were previously inaccessible due to the current limitations of quantum simulators.
@@ 2 Causality from the loop-tree duality
Generic loop integrals and scattering amplitudes, with $n$ propagators and $P$ external legs are defined in the Feynman representation as integrals in the Minkowski space of $L$ loop momenta
$\mathcal{A}_{\text{F}}^{(L)}=\int_{\ell_{1}\ldots\ell_{L}}{\cal N}\big{(}\{\ell_{s}\}_{L},\{p_{j}\}_{P}\big{)}\prod_{i=1}^{n}G_{\text{F}}(q_{i})~{},$ (1)
where the momenta of the Feynman propagators, $q_{i}^{\mu}$, are linear combinations of the loop and external momenta, and the numerator ${\cal N}$ is given by the specific theory. The integration measure in dimensional regularisation [80, 81] reads $\int_{\ell_{s}}=-\imath\mu^{4-d}\int{\rm d}^{d}\ell_{s}/(2\pi)^{d}$, with $d$ the number of space-time dimensions and $\mu$ an arbitrary energy scale. Feynman propagators are conveniently written as
$G_{\text{F}}(q_{i})=\frac{1}{\big{(}q_{i,0}-q_{i,0}^{(+)}\big{)}\big{(}q_{i,0}+q_{i,0}^{(+)}\big{)}}~{},$ (2)
where $q_{i,0}^{(+)}=\sqrt{\mathbf{q}_{i}^{2}+m_{i}^{2}-\imath 0}$ are the on-shell energies, with $q_{i,0}$ and $\mathbf{q}_{i}$ the temporal and spacial components of $q_{i}$, respectively, and $m_{i}$ the mass of the propagating particle. The infinitesimal factor $\imath 0$ is the usual complex prescription of a Feynman propagator.
Singularities in the integrand of Eq. (1) arise when Feynman propagators are set on shell, explicitly, when the energy component $q_{i,0}$ takes one of the values ±$\pm q_{i,0}^{(+)}$ in Eq. (2). An important fact to take into account is that not all potential singular configurations of the integrand in Eq. (1) lead to physical singularities of the integral. In the case of the Feynman representation of loop integrals, noncausal singularities are unavoidable. Regarding the LTD framework, its more relevant feature is the existence of a manifestly causal representation, where noncausal singularities are absent. The direct LTD representation of Eq. (1) is computed by applying the Cauchy’s residue theorem through the evaluation of nested residues [64]. To obtain the LTD causal representation we sum over all the nested residues, explicitly cancelling all the noncausal contributions. Furthermore, the LTD causal representation is found to be reinterpreted in terms of entangled causal thresholds [66, 68]. The analytical causal reconstruction is achieved by matching all combinations of $n-L$ threshold configurations that are causally compatible to each other, leading to the LTD causal representation,
$\mathcal{A}_{\text{D}}^{(L)}=\int_{\boldsymbol{\ell}_{1}\ldots\boldsymbol{\ell}_{L}}\frac{1}{x_{n}}\sum_{\sigma\in\Sigma}\frac{{\cal N}_{\sigma(i_{1},\ldots,i_{n-L})}}{\lambda_{\sigma(i_{1})}^{h_{\sigma(i_{1})}}\cdots\lambda_{\sigma(i_{n-L})}^{h_{\sigma(i_{n-L})}}}+(\lambda_{p}^{+}\leftrightarrow\lambda_{p}^{-})~{},$ (3)
with $x_{n}=\prod_{i=1}^{n}2q_{i,0}^{(+)}$ and $h_{\sigma(i)}=\pm$. The Feynman propagators from Eq. (1) are substituted in Eq. (3) by causal propagators of the form $1/\lambda_{p}^{\pm}$, with
@@
where $p$ is a partition of the on-shell energies, and $k_{p,0}$ is a linear combination of the external-momenta energy components. Given the sign of $k_{p,0}$, either $\lambda_{p}^{-}$ or $\lambda_{p}^{+}$ becomes singular after all propagators in the partition $p$ are set on shell. The combinations of entangled causal propagators represent causal threshold configurations that can occur simultaneously which are collected in the set $\Sigma$.
A crucial step in the procedure for obtaining the causal representation shown in Eq. (3), is the identification of diagram configurations admitting compatibility among causal thresholds (causal singular configurations), which is seen as querying multiple solutions over an unstructured database [77, 78]. This exercise is equivalent to select all the DAG configurations of an undirected graph in graph theory. In the following section we present a systematic and efficient quantum algorithm to address the identification of causal configurations of multiloop Feynman diagrams.
@@ 3 Quantum causal querying
The proposed algorithm recovers the central ideas from Ref. [34] on amplitude amplification or Grover’s quantum algorithm and considers a generalised description of multiloop topologies by treating them as directed graphs made of edges and eloops [70, 69]. An edge is defined as the union of multiple Feynman propagators connecting two interaction vertices. Once propagators are substituted by edges, the remaining loops of the graph define the number of eloops, which is smaller than the number of loops of the original Feynman diagram. The convenience of working with eloops in a causality scenario is given by the fact that the only feasible causal configurations are those for which the momentum flows of all propagators belonging to each edge are aligned in the same direction.
The distinguishing idea introduced in this paper is a change of logic in the construction of the oracle operator. The tagging of the unidirectional cycles representing eloops is done through the application of multicontrolled Toffoli (MCX) gates and NOT (Pauli-$X$) gates, having as a core difference with Ref. [34] the absence of binary clauses comparing adjacent edges. This logical procedure significantly simplifies the design and encoding of the corresponding quantum circuits. To evaluate the efficiency of the proposal, we focus on the number of qubits required for the algorithm implementation and on the quantum circuit depth circuit area. Before describing the algorithm, we review some relevant concepts for a better understanding of the work developed.
Regarding the efficiency of the amplitude amplification, it is important to note the relevance of the mixing angle, defined as $\theta=\arcsin\sqrt{r/N}$, that measures the ratio of the $r$ states to be queried over the total number of possible states, $N=2^{n}$, and the optimal number $t$ of the algorithm iterations. If $r\sim N/4$, then $\theta\sim\pi/6$, and one iteration is sufficient because $\sin^{2}\theta_{t}\sim 1$ with $\theta_{t}=(2t+1)\,\theta$. It has been shown from classical [66, 68] and quantum [34] analysis of the multiloop topologies depicted in Fig. 1 that the number of causal configurations is typically about one half of the total number of possible states. This implies $\theta\sim\pi/4$, which does not lead to any reasonable amplitude amplification. However, given a causal solution of a loop Feynman diagram, the mirror state with all the momentum flows reversed is also a causal solution. Therefore, for one of the qubits representing an edge, we tag only one of its states so that the number of states to be queried is $r\sim N/4$, which is the optimal value for amplitude amplification in a single iteration. The total set of causal solutions is then obtained by reversing the momentum flows of all the causal states queried. In the case that halving the number of solutions is not enough to adjust the ratio $r/N$ in the required range, then we introduce an ancillary qubit [82] to increase the number of total states without increasing the number of solutions.
@@
[…]
@@
[…]
Following to the construction of the oracle operator, it is important to recall that causal configurations are identified as DAG configurations, therefore, the eloop clauses are established to suppress directed cyclic configurations. The eloop clauses are implemented through the application of multicontrolled Toffoli gates and NOT gates that validate whether a configuration is cyclic or acyclic. To provide a convenient notation, we define the multicontrolled Toffoli operation based on Boolean clauses by
@@
where the capital $S_{j}$ contains the edge set generating the cyclic configuration involving the sub-eloop $j$. Both conditions in Eq. (6) select the same sub-eloop but with opposite edge directions. The composite condition in Eq. (7) embraces both directions of the given sub-eloop associated to $S_{j}$. Let us recall that the eloop clause register $|{a}\rangle$ is initialised to $|{1}\rangle$, therefore, applying one of the conditions in Eqs. (6) and (7) unmarks a noncausal configuration from $|{a}\rangle$.
An important aspect to consider relating eloop clauses is whether or not the set $S_{j}$ involves a fixed qubit, that is, a qubit tagged in a definite state $|{0}\rangle$ or $|{1}\rangle$. If the eloop clause condition does not involve a fixed qubit, we apply Eq. (7), which considers two multicontrolled Toffoli gates to validate both directions. The validation of both directions is stored in a single qubit $|{a_{i}}\rangle$, due to the fact that both configurations are mutually exclusive, i.e., applying one multi-controlled Toffoli gate does not overlap over the second one. In the case where the eloop clause involves a fixed qubit, the condition to probe involves only one direction, explicitly, only one of the two Boolean conditions given in Eq. (6) is tested.
Figure 2: Diffusion operator for four qubits.
@@
$U_{w}|{e}\rangle|{a}\rangle|{\text{out}}\rangle=(-1)^{f(a,e)}|{e}\rangle|{a}\rangle|{\text{out}}\rangle~{},$ (8)
[…]
$U_{s}\left(U_{w}|{e}\rangle\right)=\cos(3\theta)|{e_{\perp}}\rangle+\sin(3\theta)|{w}\rangle~{}.$ (9)
The probabilities of the orthogonal and winning states are $\cos^{2}(3\theta)$ and $\sin^{2}(3\theta)$, respectively. If $\sin^{2}(3\theta)\sim 1$, the probability of the orthogonal states is very suppressed and the measurement of the qubits in the register $|{e}\rangle$ selects the desired states with higher probabilities. The diffusion operator implemented in the quantum circuit is shown in Fig. 2 and is based on the diffusion operator described in the IBM Qiskit website  [^1] http://qiskit.org/.
The quantum algorithm presented in this paper has the advantage that the quantum circuit depth depends just on the number of eloop clauses and not on the number of edges per set, whereas the quantum circuit depth of the quantum algorithm using binary clauses [34] increases when additional edges per set are considered.
Let us recall that the quantum circuit depth [83] is defined as the integer number of gates that need to be executed along the longest path of the circuit from input to output, moving forward in time along qubit wires. The input is understood as the preparation of the qubits and the output as the measurement gate. The quantum circuit depth is computed through the Qiskit routine, QuantumCircuit.depth()  [^2] https://qiskit-test.readthedocs.io. An important feature of the quantum circuit depth is the impact of quantum gates acting on no common qubits, allowing to perform them at the same time step. On the other hand, if the gates act on at least one common qubit they have to be applied in different time-steps, increasing by one unit the quantum circuit depth.
The output of the quantum algorithm after a measurement is the selection of a single configuration. In order to identify all the solutions it is required to prepare and measure a certain number of times (shoots), having as an output a frequency histogram. An estimation of the minimal number of shots required to distinguish causal from noncausal configurations with a statistical significance of $\Delta\sigma$ standard deviations in a quantum simulator [34] is given by
$N_{\rm shots}\approx r\left(\Delta\sigma\right)^{2}\left(1+{\cal O}(\cos^{2}(\theta_{t}))\right)\,,$ (10)
assuming an efficient amplification of the causal states, i.e. $\cos(\theta_{t})\sim 0$.
@@
[…]
@@
[…]
In the following sections we explore in detail the performance of the proposed quantum algorithm by analysing the number of qubits needed for the implementation and the corresponding theoretical quantum circuit depth.
Fig. eloops (edges) Qubits $\begin{subarray}{c}\text{Quantum}\\ \text{Depth}\end{subarray}$ $\theta$ $\sin^{2}(\theta_{t})$ $\begin{subarray}{c}\text{Causal}\\ \text{states}\end{subarray}$ $\begin{subarray}{c}\text{Total}\\ \text{states}\end{subarray}$
3 two ($6$) $\phantom{0}10~{}|~{}16$ $14~{}|~{}22$ $36.8^{\circ}$ $0.87$ $23$ $64$
Table 1: Quantum circuit requirements of the quantum algorithms used to analyse the two-eloop topology in Fig. 3. The first number in the second and third columns is from the quantum algorithm proposed in this work, whereas the second number corresponds to the binary-clause algorithm [34]. The columns with a single entry have the same value for both algorithms. The numbers in the last two columns are the number of causal configurations and the total number of states, respectively.